solve linear system
3x-y=-2
5x+2y=15

Answer by checkley77(12844)

3x-y=-2 multiply by 2 & add.
5x+2y=15
6x-2y=-4
-------------
11x=11
x=11/11
x=1 ans.
3*1-y=-2
3-y=-2
-y=-2-3
-y=-5
y=5 ans.
Proof:
5*1+2*5=15
5+10=15
15=15